### v1

@@ 3 Quasi-orthogonality
and $\tilde{u}=\hat{u}$ due to $\operatorname{dist}_{a}(U,V)\leqslant\varepsilon<1$. Hence, $\mathcal{P}_{V}|_{U}$ is an injection and then is isomorphism from $U$ to $V$ since $\operatorname{dim}(V)=\operatorname{dim}(U)$.
Set $v_{j}=\frac{\mathcal{P}_{V}u_{j}}{\left\|\mathcal{P}_{V}u_{j}\right\|_{a}}$ for $j=1,2,\ldots,n$. Since $\mathcal{P}_{V}$ is an isomorphism, we have $V=\operatorname{span}(\{v_{j}\}_{j=1}^{n})$. It shows that
$\mathrel{\phantom{=}}\|u_{j}-v_{j}\|_{a}=\sqrt{\|u_{j}-\mathcal{P}_{V}u_{j}\|_{a}^{2}+\left\|\mathcal{P}_{V}u_{j}-\frac{\mathcal{P}_{V}u_{j}}{\|\mathcal{P}_{V}u_{j}\|_{a}}\right\|_{a}^{2}}$
$\leqslant\sqrt{\operatorname{dist}_{a}^{2}(U,V)+\left(1-\sqrt{1-\operatorname{dist}_{a}^{2}(U,V)}\right)^{2}}\leqslant\sqrt{2-2\sqrt{1-\varepsilon^{2}}}.$
In our analysis, we need to use the inequality $\sqrt{2-2\sqrt{1-\varepsilon^{2}}}\leqslant\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})},$ which requires that $\varepsilon\in(0,L(\theta,n))$. Here
$L(\theta,n)=\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\sqrt{1-\frac{1}{4}\left(\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\right)^{2}}.$
We see that $L(\theta,n)\in(0,\frac{1}{2})$ since $\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\in(0,\frac{1}{2})$ as $n\geqslant 2$. Consequently, we arrive at the following proposition from Corollary 3.3, which will play a crucial role in our analysis.
@@
Given $\theta\in(0,1)$ and $\varepsilon\in(0,L(\theta,n))$, if $U,V\subset H$ satisfy
$\operatorname{dim}(U)=\operatorname{dim}(V)=n,\quad\operatorname{dist}_{a}(U,V)\leqslant\varepsilon,$
then for an orthogonal basis $\{u_{j}\}_{j=1}^{n}$ of $U$, there exists $\{w_{j}\}_{j=1}^{n}\subset V$ satisfying
$a(u_{i},u_{j})$ $=a(w_{i},w_{j})=\delta_{ij},\quad i,j=1,2,\ldots,n,$
@@
$\operatorname{dist}_{a}(u_{i},w_{i})$ $\leqslant\left(1+\frac{2}{\theta^{n-3}}\left(\left(1+\theta\right)^{n-1}-\theta^{n-1}\right)\right)\sqrt{2-2\sqrt{1-\varepsilon^{2}}},\quad i=1,2,\ldots,n.$
Consider subspaces $X,Y\subset V^{h}\subset H$ with decompositions as follows:
$V^{h}=\bigoplus_{i=1}^{p}X_{i},\quad X=\bigoplus_{i=1}^{q}X_{i},\quad Y=\sum_{i=1}^{q}Y_{i},$
where $\operatorname{dim}(X)=N$ and $\operatorname{dim}(X_{i})=\operatorname{dim}(Y_{i})=d_{i}$. The following conclusion will be used in our analysis.
@@
Given $\theta\in(0,1)$, let $\varepsilon\in(0,\min_{1\leqslant i\leqslant q}L(\theta,d_{i}))$ satisfying
(8) $\max_{1\leqslant i\leqslant q}\left(1+\frac{2}{\theta^{d_{i}-3}}\left(\left(1+\theta\right)^{d_{i}-1}-\theta^{d_{i}-1}\right)\right)\sqrt{2-2\sqrt{1-\varepsilon^{2}}}<\frac{1}{\sqrt{N}}.$
If $\max_{i=1,\ldots,q}\operatorname{dist}_{a}(X_{i},Y_{i})<\varepsilon$, then
(9) $Y=\bigoplus_{i=1}^{q}Y_{i}.$
Proof 3.6.
Let $\{x_{ij}\}_{j=1}^{d_{i}}$ be an orthonormal basis of $X_{i}$ with $a(x_{ij},x_{kl})=\delta_{ik}\delta_{jl}$. We obtain from Proposition 3.4 that there exists an orthonormal basis $\{y_{ij}\}_{j=1}^{d_{i}}$ of $Y_{i}(i=1,\ldots,q)$ satisfying that for $j=1,\ldots,d_{i}$

### v2

@@ 3 Quasi-orthogonality
and $\tilde{u}=\hat{u}$ due to $\operatorname{dist}(U,V)\leqslant\varepsilon<1$. Hence, $\mathcal{P}_{V}|_{U}$ is an injection and then is isomorphism from $U$ to $V$ since $\operatorname{dim}(V)=\operatorname{dim}(U)$.
Next we set $v_{j}=\frac{\mathcal{P}_{V}u_{j}}{\left\|\mathcal{P}_{V}u_{j}\right\|}$ for $j=1,2,\ldots,n$. Since $\mathcal{P}_{V}$ is an isomorphism, we have that $V=\operatorname{span}(\{v_{j}\}_{j=1}^{n})$ and
$\mathrel{\phantom{=}}\|u_{j}-v_{j}\|=\sqrt{\|u_{j}-\mathcal{P}_{V}u_{j}\|^{2}+\left\|\mathcal{P}_{V}u_{j}-\frac{\mathcal{P}_{V}u_{j}}{\|\mathcal{P}_{V}u_{j}\|}\right\|^{2}}$
$\leqslant\sqrt{\operatorname{dist}^{2}(U,V)+\left(1-\sqrt{1-\operatorname{dist}^{2}(U,V)}\right)^{2}}\leqslant\sqrt{2-2\sqrt{1-\varepsilon^{2}}},$
i.e., $\{v_{j}\}_{j=1}^{n}$ is $\sqrt{2-2\sqrt{1-\varepsilon^{2}}}$-quasi-orthogonal. Then by Corollary 3.3, we complete the proof.
Dimension-preserving
For $V^{h}=\bigoplus_{i=1}^{p}X_{i}$, we consider subspaces $X,Y\subset V^{h}\subset H$ with decompositions as follows:
$X=\bigoplus_{i=1}^{q}X_{i},\quad Y=\sum_{i=1}^{q}Y_{i},$
where $\operatorname{dim}(X)=N$ and $\operatorname{dim}(X_{i})=\operatorname{dim}(Y_{i})=d_{i}$. Obviously, $\operatorname{dim}(Y)\leqslant\sum_{i=1}^{q}\operatorname{dim}(Y_{i})=N$. We have the following conclusion telling when $\operatorname{dim}(Y)=N$ holds and will be used in our analysis. The proof is given in Section A.2.
@@
Given $\varepsilon\in(0,\min_{1\leqslant i\leqslant q}\sqrt{\frac{4(1+\sqrt{d_{i}})^{2}N-1}{4(1+\sqrt{d_{i}})^{4}N^{2}}})$. If
$\max_{i=1,\ldots,q}\operatorname{dist}(X_{i},Y_{i})<\varepsilon,$
then
(5) $Y=\bigoplus_{i=1}^{q}Y_{i}.$
An application of quasi-orthogonality
Note that Proposition 2.3 tells only that each orthonormal eigenfunction of Eq. 3 approximates some eigenfunction of Eq. 1. However, these exact eigenfunctions of Eq. 1 may not be orthonormal to each other (see Corollary 2.11 in [11]). In practical applications, the approximate property of the approximate eigenfunctions to the exact orthonormal eigenfunctions is usually required, which is for structure-preserving and preventing the accumulation of errors of approximations.
From solving Eq. 3, we may obtain orthogonal or quasi-orthogonal (when the algebraic error of the orthogonalization is taken into account) normalized basis $\{u_{ij}^{h}\}$ of $M_{h}(\lambda_{i})$. With the investigation of the quasi-orthogonality, we are able to estimate the approximation error of the approximate eigenfunctions to the exact orthonormal eigenfunctions.
Theorem 3.7.
Let $\{u_{ij}^{h}\}$ be solutions of Eq. 3. If $\operatorname{dist}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\leqslant\epsilon_{*}$, then there exists an orthonormal basis $\{u^{o}_{ij}\}$ of $M(\lambda_{i})$ with $b(u^{o}_{ij},u^{o}_{kl})=\delta_{ik}\delta_{jl}$ such that
@@
$\operatorname{dist}\left(u^{o}_{ij},u^{h}_{ij}\right)\leqslant C_{**}\operatorname{dist}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right),\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $C_{**}=\max_{i=1,\cdots,q}\frac{2(\sqrt{d_{i}}+1)}{\sqrt{\lambda}_{i}}C_{*}$. The constants $\epsilon_{*}$ and $C_{*}$ are defined in Proposition 2.3, and $C_{*}$ is independent of $V^{h}$.
Proof 3.8.
For the approximate eigenpairs $\left\{(\lambda^{h}_{ij},u^{h}_{ij})\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}$, we obtain from Proposition 2.3 that there exist $\hat{u}_{ij}\in M(\lambda_{i})$ and the constant $C_{*}$ that is defined in Proposition 2.3 and is independent of $V^{h}$, such that
$\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|\leqslant C_{*}\operatorname{dist}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right),\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
which together with the fact $\|u_{ij}^{h}\|_{a}=\sqrt{\lambda_{ij}^{h}}$ and $\lambda_{ij}^{h}\geqslant\lambda_{i}$ yields that
$\left\|\frac{u^{h}_{ij}}{\|u^{h}_{ij}\|}-\frac{\hat{u}_{ij}}{\|\hat{u}_{ij}\|}\right\|=\frac{1}{\sqrt{\lambda^{h}_{ij}}}\left\|u^{h}_{ij}-\hat{u}_{ij}+\left(1-\frac{\|u^{h}_{ij}\|}{\|\hat{u}_{ij}\|}\right)\hat{u}_{ij}\right\|$
$\leqslant$ $\frac{1}{\sqrt{\lambda}_{i}}\left(\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|+\left|\|\hat{u}_{ij}\|-\|u^{h}_{ij}\|\right|\right)\leqslant\frac{2}{\sqrt{\lambda}_{i}}C_{*}\operatorname{dist}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right).$
Then $\{\frac{\hat{u}_{ij}}{\|\hat{u}_{ij}\|}\}_{j=1}^{d_{i}}$ is $\frac{2}{\sqrt{\lambda}_{i}}C_{*}\operatorname{dist}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)$-quasi-orthogonal and it follows from Corollary 3.3 that there exists an orthonormal basis $\{u^{o}_{ij}\}$ of $M(\lambda_{i})$ with $b(u^{o}_{ij},u^{o}_{ik})=\delta_{jk}$ such that
$\operatorname{dist}\left(u^{o}_{ij},u^{h}_{ij}\right)\leqslant C_{**}\operatorname{dist}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right),\quad j=1,\ldots,d_{i},$
where $C_{**}=\max_{i=1,\cdots,q}\frac{2(\sqrt{d_{i}}+1)}{\sqrt{\lambda}_{i}}C_{*}$. We complete the proof.
We see that Theorem 3.7 tells that, for the finite dimensional approximation of an eigenvalue problem, there exists a set of orthonormal eigenfunctions whose distance to the orthonormal approximate eigenfunctions is controlled by the distance of subspaces.
In the next section, we will present the approximation of iterative solutions $\left\{\left(\lambda_{ij}^{(n)},u_{ij}^{(n)}\right)\right\}$ produced by ParO to the solutions of the discrete problem Eq. 3 using Theorem 3.7. Then we obtain the approximation errors of iterative solutions to solutions of Eq. 1 from Proposition 2.2, Theorem 3.7 and the triangle inequalities
$\operatorname{dist}(u_{ij}^{o},u_{ij}^{(n)})$ $\leqslant\operatorname{dist}(u_{ij}^{o},u^{h,o}_{ij})+\operatorname{dist}(u^{h,o}_{ij},u_{ij}^{(n)}),$
4 Convergence and Error Estimates
With the quasi-orthogonality, in this section, we are able to obtain the convergence and error estimates of the numerical approximations produced by ParO for clustered eigenvalue problems.
4.1 Algorithm framework
We first recall the framework of ParO for the first $N$ clustered eigenvalues and their corresponding eigenfunctions of Eq. 1, which is stated as Algorithm 1. We mention that Algorithm 1 is indeed a modified version of Algorithm 1.1 in [10] (see Fig. 1 for its flowchart). We see form Fig. 1 clearly that our framework has a feature of two level parallelization in Step 2: one level is obtained by updating $N$ eigenfunctions intrinsically in parallel; the other level is obtained by applying the standard parallel strategies when updating each eigenfunction.
Algorithm 1 A framework for ParO
  1. 1.
    Given a finite dimensional subspace $V^{h}$ and initial data $\left(\lambda_{k}^{(0)},u_{k}^{(0)}\right)\in\mathbb{R}\times V^{h}$ for $k=1,2,\ldots,N$, let $n=0$.
  2. 2.
    For $k=1,2,\ldots,N$, update $u_{k}^{(n)}$ to obtain $u_{k}^{(n+1/2)}$ in parallel.
  3. 3.
    Let $u_{k}^{(n+1)}=u_{k}^{(n+1/2)}$ for $k=1,\ldots,N$. Or if necessary, project Eq. 1 onto $U_{n+1}=\operatorname{span}\left\{u_{1}^{(n+1/2)},\ldots,u_{N}^{(n+1/2)}\right\}$ and obtain eigenpairs $\left(\lambda_{k}^{(n+1)},u_{k}^{(n+1)}\right)$.
  4. 4.
    If not converge, let $n=n+1$ and go to Step $2$.
Figure 1: Flowchart of Algorithm 1
Step 1
  • physical observation or data (see, e.g., [10, 26]). For instance, in electronic structure calculations, we may choose initial data from Gaussian-type orbital, Slater-type orbital, atomic orbital based guesses, and so on;
  • solving the eigenvalue problem on a coarse grid;
  • neural networks based guesses [15, 34].
Since we look for clustered eigenvalues and their corresponding eigenfunctions, we shall consider the approximation of each eigenspace as mentioned in Introduction. We understand that it is not trivial to obtain the multiplicity $d_{i}$ of each eigenvalue $\lambda_{i}$. A possible way to approximate the multiplicities is to cluster the initial guesses $\lambda_{1}^{(0)}\leqslant\lambda_{2}^{(0)}\leqslant\cdots\leqslant\lambda_{N}^{(0)}$. By clustering methods such as Bayesian Information Criterion and Silhouette method (see e.g., [29, 33]), we can get $q^{\prime}$ clusters with $d_{i}^{\prime}$ eigenpairs in the $i-$th cluster $(i=1,\ldots,q^{\prime})$, that is,
In this paper, we assume that $q^{\prime}=q$ and $d_{i}^{\prime}=d_{i}$ for $i=1,\ldots,q^{\prime}$. Such an assumption is likely to hold since the previously mentioned methods may be able to give good initial data. We also want to point out that, the numerical experiments in [10, 24] have shown that such an assumption is unnecessary in practical computations.
Set
then $U_{n}=\sum_{i=1}^{q}U_{n}^{(i)}$.
Step 2
To update each eigenfunction in step $2$ of Algorithm 1, as pointed out in [10], we can apply the shifted-inverse approach, Chebyshev filtering, and so on. Define $\mathcal{F}_{n}:U_{n}\rightarrow U_{n+1}$ as follows
$\mathcal{F}_{n}^{(i)}:=\mathcal{F}_{n}|_{U_{n}^{(i)}}:u_{ij}^{(n)}=u_{\sum_{r=0}^{i-1}d_{r}+j}^{(n)}\mapsto u_{\sum_{r=0}^{i-1}d_{r}+j}^{(n+1/2)}:=u_{ij}^{(n+1/2)},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}).$
Set
We have $U_{n+1}=\sum_{i=1}^{q}U_{n+1/2}^{(i)}$.
Step 3
In step $3$ of Algorithm 1, if we choose to update the approximations, we can solve a small scale eigenvalue problem as follows: find $(\lambda^{(n+1)},u^{(n+1)})\in\mathbb{R}\times U_{n+1}$ satisfying
to obtain eigenpairs $(\lambda_{ij}^{(n+1)},u_{ij}^{(n+1)})$ with $b\left(u_{ij}^{(n+1)},u_{kl}^{(n+1)}\right)=\delta_{ik}\delta_{jl}$ for $(1,1)\leqslant(i,j),(k,l)\leqslant(q,d_{q})$.
Error estimate
The following theorem shows the approximation errors of eigenpairs when solving a small scale eigenvalue problem is carried out under the assumption that eigenfunctions are approximated well to a certain in Step 2. The proof is given in Section A.3.
Theorem 4.1.
If $\operatorname{dist}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)})\ll 1(i=1,\ldots,q)$, then after solving a small scale eigenvalue problem Eq. 8 in $U_{n+1}=\sum_{i=1}^{q}U_{n+1/2}^{(i)}$, there exists an orthonormal basis $\{u^{h,o}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $b(u^{h,o}_{ij},u^{h,o}_{kl})=\delta_{ik}\delta_{jl}$ such that for $(i,j)\leqslant(q,d_{q})$
$|\lambda_{ij}^{(n+1)}-\lambda^{h}_{ij}|\leqslant$ $\lambda_{i+1,1}^{h}\operatorname{dist}^{2}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\leqslant\lambda_{i+1,1}^{h}\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}^{2}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}),$
$\operatorname{dist}(u^{h,o}_{ij},u_{ij}^{(n+1)})\leqslant$ $\tilde{C}_{**}\operatorname{dist}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\leqslant\tilde{C}_{**}\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}),$
where the constant $\tilde{C}_{**}$ is independent of $U_{n+1/2}$.
We will see in the next subsection that the requirements $\operatorname{dist}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)})\ll 1(i=1,\ldots,q)$ can be achieved. We understand that the generation and solution of the $N$-dimensional eigenvalue problem in Step $3$ of Algorithm 1 is computationally expensive. Fortunately, we see from the numerical experiments in [10, 24] that the resulting matrix is almost diagonal: the non-diagonal entries are very small and the computational cost is not very high.
4.2 Shifted-inverse based ParO algorithm
In this subsection, we shall study the ParO algorithm for solving the clustered eigenvalue problem when the shifted-inverse approach is applied to update each eigenfunction. The shifted-inverse based ParO algorithm (Algorithm 3), which solves a small scale eigenvalue problem in each iteration to update the shifts, has been proposed in [10, 24]. To show the convergence, we first consider a simplified version (Algorithm 2) which fixes the shifts and does not carry out the steps of solving projected eigenvalue problems in iterations. Based on the numerical analysis of the simplified version, we then prove that the approximations produced by Algorithm 3 converge rapidly.
Simplified shifted-inverse based ParO algorithm
We set the shift as any convex combination of $\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}$ denoted by
In our discussion, we assume that the shifts are always not equal to any eigenvalue of Eq. 3 in the calculation process. Otherwise, we continue the iterative process on other eigenfunctions while keeping the eigenfunctions unchanged.
The shifted-inverse approach $\mathcal{F}_{n}$ writes: for $U_{n}^{(i)}=\operatorname{span}\left\{u_{i1}^{(n)},\ldots,u_{id_{i}}^{(n)}\right\},$ $\mathcal{F}_{n}^{(i)}:=\mathcal{F}_{n}|_{U_{n}^{(i)}}:U_{n}^{(i)}\rightarrow U_{n+1/2}^{(i)}$ with $u_{ij}^{(n+1/2)}=\mathcal{F}_{n}^{(i)}u_{ij}^{(n)}$ for $i=1,\ldots,q,j=1,2,\ldots,d_{i}$ satisfying
The simplified shifted-inverse based ParO algorithm is stated as Algorithm 2. Compared with Algorithm 1, step 3 is no longer carried out in this simplified version.
Algorithm 2 Simplified shifted-inverse based ParO algorithm
  1. 1.
    Given a finite dimensional space $V^{h}$ and $tol>0$, provide and cluster initial data by Eq. 6, i.e., $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}\subset\mathbb{R}\times V^{h}$. Set $\bar{\lambda}_{i}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)$ and let $n=0$.
  2. 2.
  3. 3.
    Set $u_{ij}^{(n+1)}=\frac{u_{ij}^{(n+1/2)}}{\left\|u_{ij}^{(n+1/2)}\right\|_{b}}$. If $\frac{\|u_{ij}^{(n+1)}-u_{ij}^{(n)}\|_{b}}{\|u_{ij}^{(n)}\|_{b}}>tol$, let $n=n+1$ and go to Step $2$.
If the initial guesses approximate the exact eigenvalues well enough, then the source problems Eq. 9 will be ill-conditioned. We mention that there are approaches to deal with ill-conditioned systems (see e.g., [4, 6, 7, 28, 30]). Indeed, it is quite difficult to solve these ill-conditioned systems well, which will be discussed in our other work. Here we assume that such systems can be well solved.
The following proposition tells that the approximation to the eigenspace in the iteration process is of dimension-preserving. The proof is given in Section A.4.
@@ 4.2 Shifted-inverse based ParO algorithm
If $U^{(i)}_{n+1/2}$ is obtained by Algorithm 2, then
With the help of Proposition 4.2, we obtain convergence of eigenspace approximations produced by Algorithm 2. The proof is given in Section A.5.
Theorem 4.3.
Assume that
(11) $\operatorname{dim}\left(U_{0}^{(i)}\right)=d_{i},\quad\operatorname{dist}\left(M_{h}(\lambda_{i}),U_{0}^{(i)}\right)<1\quad\forall i=1,2,\ldots,q.$
If $U_{n+1/2}^{(i)}$ is produced by Algorithm 2, then